Find I, in the circuit Ix- 10 mA I, 2 mA 2. Determine Vo- & VI 12 kI
Answer #1

Only one question can be posted as per HOMEWORKLIB RULES

1.

From Kirchoff's Junction rule

-Ix-10+5Ix+2=0 (Incoming current to node taken as positive and outgoing current taken as positive)

4Ix=8

Ix =8/4 =2 mA